OpenStudy (anonymous):

perform the operations and simplify (x+y)-2x(x-y)/(x+y)(x-y)

OpenStudy (anonymous):

((x+y)/(x+y)(x-y)) - (2x(x-y)/(x+y)(x-y)) (1/(x-y)) - (2x/(x+y)) I think this is simple enough, isn't it?

OpenStudy (anonymous):

think we h ave a common factor top and bottom yes?

OpenStudy (anonymous):

It does not make since

OpenStudy (anonymous):

or is it : (x+y) - (2x(x-y)/(x+y)(x-y)) ((x+y) - (2x/(x+y))

OpenStudy (anonymous):

\[\frac{(x+y)-2x(x-y)}{(x+y)(x-y)}\]

OpenStudy (anonymous):

is that it?

OpenStudy (anonymous):

It is (x+y)-2x+(x-y)/(x+y)(x-y). I am stuck on the numerator

OpenStudy (anonymous):

multiply out and get \[\frac{x+y-2x^2+2xy}{(x+y)(x-y)}\]

OpenStudy (anonymous):

why multiply when this is adding and subtracting

OpenStudy (anonymous):

or just leave it. there is nothing more you can do with it

OpenStudy (anonymous):

maybe i am not sure what the question is. is it the one i posted?

OpenStudy (anonymous):

I need to simplify this equation and I am not sure if it is simplified the way it is

OpenStudy (anonymous):

well \[\frac{(x+y)-2x(x-y)}{(x+y)(x-y)}=\frac{x+y-2x^2+2xy}{(x+y)(x-y)}\] but there is really nothing more you can do with it. nothing factors, nothing cancels

OpenStudy (anonymous):

Does it make a difference that there should be a + sign between the -2x and the (x-y)